### v1

# Combining Tests and Proofs for Better Software Verification

Abstract.
Test or prove? These two approaches to software verification have long been presented as opposites. One is dynamic, the other static: a test executes the program, a proof only analyzes the program text. A different perspective is emerging, in which testing and proving are complementary rather than competing techniques for producing software of verified quality.
Work performed over the past few years and reviewed here develops this complementarity by taking advantage of Design by Contract, as available in Eiffel, and exploiting a feature of modern program-proving tools based on “Satisfiability Modulo Theories” (SMT): counterexample generation. A counterexample is an input combination that makes the program fail. If we are trying to prove a program correct, we hope not to find any. One can, however, apply counterexample generation to incorrect programs, as a tool for automatic test generation. We can also introduce faults into a correct program and turn the counterexamples into an automatically generated regression test suite with full coverage. Additionally, we can use these mechanisms to help produce program fixes for incorrect programs, with a guarantee that the fixes are correct. All three applications, leveraging on the mechanisms of Eiffel and Design by Contract, hold significant promise to address some of the challenges of program testing, software maintenance and Automatic Program Repair.
Seeding contradiction, Program proofs, Software verification, Contracts
copyright: acmlicensedjournalyear: 2025isbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Error handling and recoveryccs: Software and its engineering Software verification and validationccs: Software and its engineering Software testing and debugging
1. From opposition to complementarity
Software development is not just an intellectual exercise but an engineering endeavor, with considerable benefits when it succeeds, but also the risk of considerable damage if the resulting software is incorrect. Software verification seeks to avoid such adverse outcomes. It has two variants, dynamic and static. Dynamic verification, also known as testing, consists of executing a system on sample input conditions and checking the results against expected properties (“oracles”). For many projects in industry, testing is the only form of verification. In contrast, “verification” among programming researchers usually refers to static methods, which do not execute the program but analyze its text, usually in its source form; the analysis may be directed towards specific properties, or it may perform a full proof of correctness, which assumes that developers have produced not only a program but also a formal specification of its intended behavior, and determines whether the program correctly implements that specification.
There is a long history of pitting tests against proofs. Practitioners often reject proofs as too hard, costly and impractical. Critics of testing, for their part, dismiss them as ineffective since they only exercise a minuscule subset of all possible cases, as famously stated by Dijkstra: “Program testing can be used to show the presence of bugs, but never to show their absence!(Dijkstra, 1972). The controversy was at times heated, with strong dismissals of proofs (Goodenough and Gerhart, 1975) (against Naur’s proof attempts (Naur, 1969)), (DeMillo et al., 1979), and equally stark statements about the limitations of testing by proponents of proofs.
In recent years, more conciliatory attitudes have developed, attested among others by the creation of the TAP conference, Tests And Proofs, running since 2007. Verification is difficult in any of its forms, and should not let any dogma exclude any potentially useful technique. Even the staunchest proof advocate must accept that it may not be possible to specify all properties formally, particular “non-functional” attributes such as performance and fault tolerance, and that the value of any proof is conditional on assumption that the compiler and hardware both preserve program semantics, leaving room for tests. One may also note the role of failures on both sides:
  • A failed test is actually a proof: it demonstrates that the program is incorrect — in the same way that in the Popperian view of science a negative experiment proves the theory wrong. (In this respect one may interpret Dijkstra’s quip, usually heard as an indictment of tests, as advertisement for them: it is one of the goals of software verification to help uncover faults.)
  • A failed proof attempt often leaves us clueless at what is wrong. To understand, it is useful to see a test triggering the underlying fault if any. Section 4 will develop this idea.
Test-and-proof complementarity goes beyond these observations. The catalyst for the work reported in this article is a feature of modern program proving tools, such as those based on SMT (Satisfiability Modulo Theories) (Barrett et al., 2010), which try to generate counterexamples violating the specification. A proof will succeed if it cannot produce any; but for an incorrect program the counterexamples yield tests. We may also use the resulting information to propose corrections to the bugs. For a correct program, we may obtain a regression test suite by injecting faults and producing counterexamples.
We have applied these ideas to develop verification and program-repair techniques that exploit the complementarity, while taking advantage of the Design by Contract approach to software construction (Meyer, 1997). Described until now in separate articles (Huang and Meyer, 2023; Huang et al., 2022, 2023, 2024a, 2024b, 2025b), these steps gain here a comprehensive and up-to-date presentation for a broader audience. Section 2 explains the general idea: taking advantage of the information associated with proof failures. Section 3 is a reminder on Design by Contract, a key ingredient of the approach. Subsequent sections introduce the successive refinements of the basic idea:
  • Proof2Test (Section 4 and (Huang and Meyer, 2023)) turns failed proofs, often seemingly unhelpful, into useful test cases, which offer concrete evidence of why the proofs fail, enabling programmers to understand and address the underlying issues.
  • Proof2Fix (Section 5 and (Huang et al., 2024a)) extends the combination of tests and proofs to automatic program repair by providing a guaranteed validation of fixes through the prover.
  • ‘Seeding Contradiction” (Section 6 and (Huang et al., 2023)), extends the capability to systematic test generation. By deliberately creating failed proofs and running them through Proof2Test, it can produce a collection of tests that satisfy code coverage criteria. It also incorporates loop unrolling to create tests that explore loop behavior and helped provide an assessment of MC/DC, a testing criterion sometimes viewed skeptically but widely used in industry.
Section 7 reviews earlier work on neighboring topics and Section 8 draws conclusions.
2. The core idea: proofs and their failures
The idea of proving programs correct (already prefigured in visionary late-1940s articles by Turing and Von Neumann) goes back to the seminal work of Floyd, Hoare, Dijkstra, McCarthy, Naur and others in the 1970s and 1980s. What has changed in recent years is the appearance of practical tools: “provers”, which can prove the correctness of programs, including large ones. Proving a significant program involves far too many steps to make human checking either practical or (if it somehow were possible) believable. The role of automation has consequently been essential. While still used by only a minority of projects, provers have achieved some industry successes (Huang et al., 2025a).
A prover works on the combination of a program and a specification (often expressed as “contracts” as explained in the next section). The starting point for the present work is to take advantage of the technique used internally by many modern provers: they seek one or more “counterexamples” — example inputs falsifying the expected properties If this attempts fails, the proof succeeds [^1] The duality between failure and success is a general theme of the present work. It follows from the generalization of De Morgan’s laws to predicate calculus: the negation of a universal is the existential of the negation, and conversely. In symbols, $\neg(\forall x\mid p)=\exists x\mid\neg p$. A proof seeks a universal ($\forall$) property; a failed test is an existential ($\exists$) property of the negation.. The present work focuses on the intermediate steps, when the proof fails, meaning the search for a counterexample succeeds; and, more specifically on the counterexamples themselves.
Successive developments show that this idea has several fruitful applications. The first is to turn counterexamples into tests. A proof failure is disconcerting: “postcondition violated” does not help the programmer understand what has been done wrong. Internally, though, the prover has a counterexample; “for input 5, the output is -3 while required to be positive” is considerably more informative for a programmer, even more so if accompanied by an automatically generated runnable test reproducing the failure. In other words, a tool can turn failed proofs into failed tests, with two major benefits: providing concrete, runnable evidence to the programmer; and yielding a regression test to be run again in the later history of the project once the bug has been fixed. Those are the functions of Proof2Test (Section 4).
The next application is to use the counterexamples to generate program fixes. Automatic program repair (APR) needs “invariants” characterizing failed executions. Instead of getting them from manually generated tests, the present approach can generate guaranteed invariants; instead of validating fixes through tests, it can guarantee their correctness by running them through the prover. Those are the functions of Proof2Fix (Section 5)
For critical systems, the test suite must reach a high level of coverage, going beyond basic “branch coverage” to such industry-standard measures as MC/DC, and unfolding loops so that the tests can exercise several successive iterations rather than just zero or one (section 6).
3. The role of contracts
While of general applicability, the results reported in this article rely on a contract mechanism as offered by various languages including the one used here, Eiffel. The need for contracts results from the observation that verifying the correctness of a program requires having a specification of its intended behavior. Contracts are elements of specification that are embedded in the code itself (rather than in separate documents) and, as a result, can be used for both testing (which will evaluate them in test runs) and proofs. They appear in several modes:
  • Simple assertions (check, in Fig. 5): properties that must hold at specific points in the code.
  • Precondition (require, lines 3–4 in Fig. 2): properties that clients must satisfy for any call.
  • Postconditions (ensure, 18 – 20 in Fig. 2): properties that the routine (supplier) guarantees.
  • Loop invariant (invariant, 9 – 11 in Fig. 2): properties guaranteed after loop initialization and every iteration.
  • Loop variant (variant, 16 in Fig. 2): integer measure that remains non-negative and decreases at each loop iteration, ensuring loop termination.
As an example, Fig. 2 shows an Eiffel class MAX. The intent of the max function is to return into Result the maximum element of an integer array a of size a.count. The postcondition clauses in lines 19 and 20 (is_max and result_in_array) specify the intent of the function: every element of the array should be less than or equal to Result; and at least one element should be equal to Result.
Contracts are a strong enabler for both testing and proving because they are specification elements embedded in the code itself. For testing, they make it possible to identify the source of a bug precisely: a precondition violation indicates a bug in the client; if the test passed the precondition, any subsequent failure is a bug in the code of the supplier method. Tools such as AutoTest (Meyer et al., 2007a) have used these ideas to provide automated random testing. For proving, most tools rely on Dijkstra’s weakest-precondition (wp) calculus (Dijkstra, 1976), which deliver their full potential when the code already includes postconditions and explicit loop variants and invariants.
Many programming languages (such as Java, C#, C++, Python) support simple assertions but have no specific constructs for the other uses listed above. Extensions to some of them have introduced such constructs: JML (Cok, 2021) for Java and Spec# for C#. A handful of languages, on the other hand, enjoy built-in contract mechanisms. They include Eiffel, SPARK, Dafny and the upcoming“C++26” ISO/IEC C++ standard. The present work is based on Eiffel because of its extensive support for contracts as a core native mechanism rather than an add-on, its practical availability through tools such as the EiffelStudio compiler and IDE, and its long experience of research on tools using contracts for automatic test generation (AutoTest), automatic program repair (AutoFix) and proofs (AutoProof, used in the present work) (Meyer et al., 2007a; Wei et al., 2010; Tschannen et al., 2015). All the tools and techniques described in this article have been integrated into the research version of EiffelStudio (Fig. 1).
Figure 1. Combination of tools for program verification, testing and fixing
Contracts are also necessary to apply the theory used by many modern provers: “Hoare logic” of which Dijkstra’s weakest-precondition (wp) calculus (Dijkstra, 1976) is a variant. The basic condition to be established for a routine $r$ of body $b$, precondition $P$ and postcondition $Q$ is
$P\implies(b\ \textbf{wp}\ Q)$
where $b$ wp $Q$ is the the weakest possible assertion such that $b$, started in a state satisfying $P$, will terminate in a state satisfying $Q$. The prover will compute this condition for every $r$ and then prove that every call satisfies it. As discussed in the next section, provers based on SMT solving achieve this goal by trying to find an assignment of variables (a counterexample) that contradicts it. The proof succeeds if they cannot. The present work exploits the converse case.
4. Creating tests from failed proofs
The observations of the previous section highlight what a failure means in the proving and testing processes, and how the two cases relate to each other. A failed proof leaves the programmer wondering what is wrong. A failed test provides concrete evidence.
For anyone trying to prove a program correct, failure is a familiar companion. Textbooks usually show successful proofs, but in practice the path to success can be frustrating. You are using a prover and prepared the next proof step diligently: the program, its specification (contracts), and extra annotations known as “verification conditions”. You click ”Verify”, and the prover tells you it is not able to perform the proof successfully. Fig. 2 shows such a response, for the AutoProof prover, before the inclusion of the tools discussed below.
Figure 2. A proof fails.
Where does one go from here? The message states that a postcondition (line 28), describing the intended effect of the routine is_max, may be violated. That does not even mean that the routine is incorrect; just that the prover cannot prove that it is correct. Is the implementation wrong (the most frequent case)? Is the specification wrong? Did the developer forget to include some verification conditions? Or is the task exceeding the prover’s abilities?
The process of getting to a successful proof has sometimes been called “static debugging”, in reference to the classical form of debugging, which is “dynamic” as it involves running the program on tests. When they succeed, static techniques provide a guarantee of correctness that no set of tests can match; but a failed test has the practical advantage of giving the programmer directly useful evidence of what is wrong, helping to identify the source of the failure — the bug (fault) — and correcting it. Such responses are known in the psychology literature as constructive feedback (Kluger and DeNisi, 1996), although the basic idea is clear from daily-life experience (“your answer is wrong” is less effective than “on page 3 you mistook pounds for kilograms”). In dealing with programming errors, the programmer needs mentally to reproduce the faulty program’s execution. A failed test gives such a directly usable indication; the failed-proof message in Fig. 2 does not.
Modern proof tools have the potential to tell us more. To establish Hoare-style properties as discussed in the previous section, AutoProof relies on Boogie (Le Goues et al., 2011) which itself uses an “SMT solver” such as Z3 (Barrett et al., 2010), whose method to prove a property consists of looking for counterexamples that defeat it. The proof succeeds when it cannot find any. It fails if it finds one or more. To gain more concrete information, the programmer could then in principle explore the solver’s internals for counterexamples. Often, however, that information is spread over hundreds of lines in internal files whose format, SMT-LIB (Barrett et al., 2010), is intended for verification experts. Tools such as BVD (Boogie Verification Debugger) and Boogaloo can help but remain largely static.
Proof2Test (Huang and Meyer, 2023) extends AutoProof and provides the missing step. It analyzes internal information to display a directly understandable input example causing the failure, consisting here of the array’s size and some of its elements, as shown in Fig.  3 (a) and then generates a test case as in Fig.  3 (b).
Figure 3. A failed proof providing a concrete (minimized) erroneous case
In Fig.  3 (a) the values are small: a two-element array with values 0 and 1. If asked to produce one counterexample, the SMT solver typically yields much larger values. Here it initially produces an array size a.count of 11,800 and values a [11799] ==\, 0 and a [11800] ==\, 5. For disproof purposes (showing that the postcondition does not hold), they are as good as any others. Since the purpose of Proof2Test is to provide concrete evidence helping programmers figure out concretely what is going on, we prefer values that mean something at the human scale. Proof2Test takes advantage of the underlying prover’s ability to generate not just one but several counterexamples (when provided with different seeds). It applies a minimization algorithm (Huang et al., 2022) to generate successive ones until it finds minimal values, such as those of Fig.  3. As with traditional programmer-devised test cases, they enable the programmer to relate to the failure by trying them out, as in Fig. 2, and to see that on loop initialization at lines 10 – 12 Result = 0 and i = 2 so that at line 13 the exit condition evaluates to True, terminating the loop and revealing the fault: early termination prevents the program from getting to the maximum value at position 2. To eliminate the error, it suffices to strengthen the exit condition to permit one more loop iteration: change i $\geq$ a.count to i >> a.count.
In addition to producing information that helps programmers understand and correct the fault behind a failure, Proof2Test can turn the counterexample into a test. Fig. 3 (b) is actually an executable test case produced by Proof2Test, expressed in the programming language. Programmers can run the test and observe, step by step, how it breaks the specification:
  • Create an instance current_object of class MAX (line 7).
  • Create an integer array a and fill it with values 0 at position 1 and 1 at index 2 (lines 8 – 9).
  • Call the erroneous function max on current_object with a as argument (line 10).
Running this code in AutoTest produces a run-time failure from the violation of the postcondition is_max, providing tangible evidence (not available from the failed proof attempt) of what is wrong with the version of Fig. 2 and making it possible to resume the development process.
A preliminary evaluation of Proof2Test on 20 Eiffel programs (Huang and Meyer, 2023) demonstrated the approach’s potential. With an Intel 12-Core processor and 32 GB of RAM, the above process takes less than 0.5 seconds on average. In most cases, the generated failing test is useful: executing it yields a specific trace illustrating how the program leads to the same contract violation that makes the proof fail; programmers can use the debugger’s step-by-step mode to understand the issue.
In a small proportion of cases, the generated test does not lead to a failure. The reason is usually that the implementation is “correct” in some intuitive sense — it does what the programmer somehow intended — but the specification is incomplete. These passing test runs are useful too, although in a different way from the failing ones: they alert the developers to a problem in the specification. (We can hardly say that a program is correct if we are not able to state precisely what it is supposed to do.) In either case, the generated tests are important as regression tests: once the bugs have been corrected, every previously failing test should become part of the project’s regression suite.
Experiment results (Huang et al., 2022) also show that minimization is cost-effective: in most cases, it reduces the values of integer variables by over 80% with an average cost of less than 4 extra reverification runs. Most minimized values are small and relatable: out of 125, 108 are in the range [-2, 2], out of which 58 are zero; others are usually close to some constants appearing the program.
5. Combining tests and proofs for better program repair
Identifying bugs is good; correcting them is better. Automatic program repair (APR) has demonstrated its potential in producing useful fixes. The Achilles’ heel of most existing APR approaches (Monperrus, 2018) is that they rely on test cases for both bug identification and fix validation. Writing test cases is tedious; running them can be time-consuming; and (the most significant limitation) validating them through tests does not guarantee their correctness (Dijkstra again). Proof2Fix (Huang et al., 2024a), based on Proof2Test, implements a static APR approach.
The tool chain (Fig. 4) is derived from the standard workflow in APR by replacing tests with proofs. The process starts with verifying a program using a prover (here AutoProof); if the prover detects a fault, Proof2Test generates counterexamples, each of which represents a failing execution trace, in which the program goes to a state that violates a desired property. Traditional APR techniques analyze failing tests (usually generated manually) to identify any common elements which might pinpoint the cause of the failure; Proof2Fix uses the same idea but with tests replaced by counterexamples generated automatically and statically.
Figure 4. Proof2Fix workflow
The generation of abstract general properties from counterexamples uses Daikon (Ernst et al., 2007), an inference tool which produces a set of counterexample invariants: predicates on the input variables which hold in the counterexamples, matched to a set of patterns (equality between variables or with constants, linear relations e1 ==\, a $\cdot$ e2 + b or e1 ==\, a $\cdot$ old e2 + b etc.) The invariants make it possible to produce candidate fixes for either the contract or the implementation. (Empirical studies (Meyer et al., 2007b) indicate that bugs arise from both kinds of mistake.) AutoProof then filters out invalid fixes, retaining only the fixes that pass the proof (they remove the original proof failure and introduce no new one). Valid choices can then be presented to the programmer for decision. Neither at the bug-finding stage nor at the bug-fixing stage is there any need to invent test data or to produce a test harness.
Evaluation of fixing 80 proof failures shows (Huang et al., 2024a) that the approach can produce meaningful fixes, formally validated. It generates at least one valid fix for 82.5% of all failures and at least one programmer-approved fix for 37.5% of them. A fixing session takes about 1 minute. Proof2Fix is most effective for failures caused by two types of faults: incorrect source expression in an assignment and incorrect condition in a conditional instruction. In both cases, the inferred counterexample invariants correctly characterize the faulty cases that need to be ruled out. The approach is, on the other hand, not good at fixing failures caused by too-weak contracts, as the generated counterexamples exhibit too much diversity.
6. Proofs as a foundation for efficient test suite generation
When the verification results in multiple failures, Proof2Test produces a number of different tests. The Seeding Contradiction strategy (SC) (Huang et al., 2023), takes advantage of this possibility.
Every serious software project needs a regression test suite, essential to managing the project’s evolution by ensuring that new developments do not invalidate previously working functionality. (Software engineering history includes numerous examples of regression bugs, such as the 2012 Knight Trading Group bug, which lost the firm \$440 million in 45 minutes and almost bankrupted it, and the July 2024 CrowdStrike Falcon Sensor update.) Aside from the practical advice of including a test for every case that failed at some point in the project’s history, it is very difficult to produce a regression test suite with a high coverage of all the possible execution paths.
At first sight, the Proof2Test techniques described above seem inapplicable, since they deal with faulty programs, whereas we need a regression test suite for a working program. The idea behind Seeding Contradiction is, as the name suggests, that we will make a correct program incorrect — in many different ways — by inserting (“seeding”) a bug in every branch. While counterintuitive at first (the usual goal in software engineering is to go from incorrect to correct!), this idea enables us to benefit from the preceding techniques, which generate counterexamples for faulty programs. Under suitable conditions, a counterexample for an incorrect program can be an example (a case of correct behavior) for the corresponding correct program.
Proof2Test, with Seeding Contradiction, inserts a faulty instruction (“trap property” (Gargantini and Heitmeyer, 1999)), into every control branch, as shown in Fig. 5. The instruction is a special case of the instruction
Figure 5. A two-branch program with seeded contradictions
check p end which states that the boolean property p must hold at the corresponding program point, causing the verification to fail if it cannot guarantee it. (Execution, if attempted, would cause an exception if p does not hold at run time.) With False for p, the branch will fail to verify and the Proof2Test mechanisms of the previous section will generate the corresponding counterexample. The technique actually works for an incorrect program as well as for a correct one. Even though it uses a prover normally intended for programs equipped with contract elements (preconditions, postconditions, class and loop invariants, loop variants), it is applicable to uncontracted programs as in this example.
By construction, the generated test suite achieves 100% branch coverage for feasible branches (excluding unreachable ones, which no test can possibly exercise). Proof2Test goes beyond branch coverage in two complementary ways:
  • MC/DC. On option, Proof2Test enforces the “Modified Condition/Decision Coverage” criterion, recommended in several important industrial software safety standards particularly observed in the aerospace industry. [^2] Examples are DO-178B (Software Considerations in Airborne Systems and Equipment Certification) from the Radio Technical Commission for Aeronautics and ISO 26262-3:2011 (Road Vehicles - Functional Safety).
  • Loop unrolling. The body of a standard “while” or “until” loop, by definition, can be executed a variable number of times; but branch coverage only needs a path with zero execution and another with at least one execution. In other words, it treats a loop like a conditional, missing bugs that arise only with other numbers of execution. Proof2Test includes a loop unrolling mechanism, which considers loop bodies executed (“unrolled”) any number of times, up to a user-settable maximum determined in consideration of performance constraints.
Our measurements of the results of enforcing MC/DC and loop unrolling (Huang et al., 2025b) indicate that both enhancements uncover bugs that plain branch coverage misses. The example programs, while significant and in some cases sophisticated (coming for example from verification competitions) are still small, so we refrain from sweeping conclusions, but the results are encouraging:
  • MC/DC increases the number of found bugs by 12.2% over branch coverage, growing to 14.6% for programs with complex multiple-condition decisions (Huang et al., 2024b). This result is significant since there is a certain skepticism towards MC/DC in the testing community (due in part to the original papers’ lack of empirical results); some authors even dismiss the approach (Bhansali, 2007).
  • Loop unrolling uncovers more faults as the depth of unrolling increases. Measurements show an improvement of 14.5% over plain branch coverage with 5-level unrolling, uncovering 86.1% of all faults (Huang et al., 2025b). The rise of detected faults is steep for small unrolling levels ($\leq$ 5); going from 1 to 2 brings a 9% improvement. Beyond 5, the effect decreases (by 3% from 5 to 8).
Many advanced test generation techniques, such as directed automated random testing and concolic testing (Sen et al., 2005), still require some code execution. The techniques described here are entirely static: they perform verification, using a prover, but do not execute the program and hence do not require test data (although they generate such data) or test harnesses. As noted, they work on correct as well as incorrect programs. If a program calls a routine with a precondition, the generated tests will satisfy the preconditions thanks to SC’s constraint-solving techniques.
The results summarized above, from empirical studies (Huang et al., 2023, 2024b, 2025b), are achieved with reasonable performance. The SC strategy produces a test suite much faster than some of the dynamic techniques, such as IntelliTest (Tillmann and De Halleux, 2008) (previously known as Pex, a symbolic execution test-generation tool for .NET) and AutoTest, a test generation tool for Eiffel using Adaptive Random Testing, specifically ARTOO (Ciupa et al., 2008). For a typical small program, SC requires approximately 0.5 seconds to generate a 100%-branch-coverage test suite; that is about 50 times faster than IntelliTest and 500 times faster than AutoTest. Adding MC/DC multiplies the generation time by about 2 and the test suite’s size (an important concern in practice since a good software process runs test suites frequently) by a factor of less than 10. Adding unrolling for test generation, depending on the complexity of the loop’s control-flow structures, would lead to different increment on generation time. For plain loops (with neither embedded conditionals nor nested loops), it will lead to a roughly linear growth of generation time as the unrolling depth increases. While for nested loops or loops with more complex conditionals embedded inside, as more failing instructions are introduced during test generation, the increment of generation time becomes more substantial. The evaluation in (Huang et al., 2025b) shows that for the programs with nested loops, the time increases exponentially and becomes unacceptably high (over an hour) when the unrolling depths reaches 8.
7. Related work
A number of previous or parallel efforts turn counterexamples, generated from failed verification, into diagnosis messages. The Boogie Verification Debugger (BVD) (Le Goues et al., 2011) allows programmers, statically, to “debug” Boogie proof-failure reports statically (Le Goues et al., 2011). Other tools (Chakarov et al., 2022; Müller and Ruskiewicz, 2011) pursue similar goals. These approaches provide static traces for analysis; the approach reported here (Section 4) goes further by producing a dynamic trace and actual tests that programmers can run to see concretely what is going on.
Many testing projects have used SMT solving techniques to generate tests. Some, such as Klee (Cadar and Nowack, 2021), PathCrawler (Williams et al., 2005), Pex/IntelliTest (Tillmann and De Halleux, 2008) and CUTE (Sen et al., 2005) are based on symbolic execution. They explore paths in a program and use a constraint solver to reason about their feasibility. These strategies include a dynamic component and as a result cannot provide guarantees of exhaustive branch coverage. The testing strategies reported in the present work (Section 6) are static.
The earliest work we know to have applied this idea (Angeletti et al., 2009) generates tests for low-level C programs using Bounded Model Checking (BMC) (Kroening and Tautschnig, 2014), producing test suites with exhaustive branch coverage. A more recent variant, for Java bytecode, is JBMC (Brenguier et al., 2023), where each verification run only activates one assertion at a time, producing one counterexample (the C version (Angeletti et al., 2009) uses compile-time macros, one for each block to avoid actually generating multiple programs); in contrast, the SC framework produces only one program, using a single run-time variable representing the block number. Similarly, (Beyer et al., 2004) extends the software model checker Blast to generate test suites of full coverage with respect to given predicates. BMC-based approaches rely on the correctness of the bound of the execution trace: if the bound is not set correctly, some branches might not be covered, requiring more verification runs to obtain a better bound. The SC approach does not need such mechanisms. An approach (Nilizadeh et al., 2022) that applies ideas of fault injection for generating tests (building on work on using counterexamples for program repair (Nilizadeh et al., 2021)) in Hoare-style verification exploits counterexamples produced by the OpenJML (Cok, 2021) verification tool to generate unit tests in JUnit format. These approaches all bear similarities to the present work, which seems to provide the most far-ranging and unifying collection of proof-test combinations techniques, from exploiting proof failures to counterexample minimization, test generation covering several code coverage measures and automatic program repair.
Automated test generation is an active research area (e.g. (Gorla et al., 2025)) which in recent years has increasingly used LLM, generally leading to average coverage levels of between 75% to 90%. Proof2Test is much more efficient. LLMs can also serve to generate specifications automatically (Ma et al., 2025), although questions remain: while such specifications might make the code provable, they might also be the wrong specifications, preventing the generation of relevant some tests.
The present work distinguishes itself by the use of Eiffel. While it can be viewed as a limitation, this choice also provides a considerable advantage since it does require an addition to another language (as JML for Java) but uses an existing (ISO-standardized) language, taking advantage of the built-in contract mechanism and hence directly using the original contract-equipped source code for both tests (with EiffelStudio’s mechanism for evaluating assertions during testing) and proofs. Users of other language may view the present results as those of “laboratory work”, using ideal conditions so as to open the way for others to extend the results to other environments.
8. Conclusions
The work presented here suffers from a number of limitations. Examples so far are still small, although some involve sophisticated algorithms. Some of the applications (but not all) assume that the programs have been equipped with contracts, a step that not all programmers are prepared to take. The program-repair applications so far have mostly covered variables of basic types and need to be extended to complex data structures.
These results, however, are promising, in particular on the performance side: verification is faster than running many tests, an advantage which grows rapidly with the program’s size and complexity. More generally, any arguments along the lines that programmers allegedly do not want to write contracts should be weighed against the tasks that programmers have to carry out in today’s dominant test-based approach. The commonly used phrase “automatic testing”, in its current meaning, is a misnomer, obscuring the unpleasant reality: the enormous burden that test case generation, still largely a manual task, imposes on programmers. (AI tools will help, but they will also help writing contracts.) With a static approach, that task can be automated.
More generally, program proofs and dynamic tests pursue the same ultimate goal of software correctness. This article has shown that one can treat them as complementary rather than exclusive. We have seen three major applications of the idea, exploiting the features of SMT solvers underlying modern proof tools: turning failed proofs, often elusive to the programmer, into directly usable tests (especially after minimization) evidencing the failure, which brings examples to human scale; generating 100%-coverage test suites entirely automatically; and providing guaranteed bug fixes. Today’s tools are still imperfect but they point the way to a major leap in the effectiveness of software verification and repair, combining the best of what the software engineering community has learned both on the side of tests and on the side of proofs.